OpenStudy (anonymous):

The sum of the interior angles of a polygon is equal to three times the sum of its exterior angles. How many sides does the polygon have?

OpenStudy (anonymous):

Do you remember the formula of the sum of the interior angles of a polygon?

OpenStudy (anonymous):

yap

OpenStudy (anonymous):

What is it?

OpenStudy (anonymous):

180(s – 2)

OpenStudy (anonymous):

Now what is the equation for the sum of the exeterior angles?

OpenStudy (anonymous):

360

OpenStudy (anonymous):

interior angles of a polygon is equal to three times the sum of its exterior angles

OpenStudy (anonymous):

\[ 180^\circ(s-2)=3\times 360^\circ \]

OpenStudy (anonymous):

Shouldn't you just solve for \(s\)?

OpenStudy (anonymous):

yap

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

is it 4

OpenStudy (anonymous):

I don't think so.

OpenStudy (anonymous):

If you divide both sides by \(180^\circ\) you get: \[ s-2=3\times 2 \]

OpenStudy (anonymous):

is it 8

OpenStudy (anonymous):

I think \(8\) is right.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Great job.

OpenStudy (anonymous):

You're smart enough to do it.